### v1

# Does a Rising Tide Lift All Boats? Bias Mitigation for AI-based CMR Segmentation

Abstract
Artificial intelligence (AI) is increasingly being used for medical imaging tasks. However, there can be biases in the resulting models, particularly when they were trained using imbalanced training datasets. One such example has been the strong race bias effect in cardiac magnetic resonance (CMR) image segmentation models. Although this phenomenon has been reported in a number of publications, little is known about the effectiveness of bias mitigation algorithms in this domain. We aim to investigate the impact of common bias mitigation methods to address bias between Black and White subjects in AI-based CMR segmentation models. Specifically, we use oversampling, importance reweighing and Group DRO as well as combinations of these techniques to mitigate the race bias. Furthermore, motivated by recent findings on the root causes of AI-based CMR segmentation bias, we evaluate the same methods using models trained and evaluated on cropped CMR images. We find that bias can be mitigated using oversampling, significantly improving performance for the underrepresented Black subjects whilst not significantly reducing the majority White subjects’ performance. Group DRO also improves performance for Black subjects but not significantly, while reweighing decreases performance for Black subjects. Using a combination of oversampling and Group DRO also improves performance for Black subjects but not significantly. Using cropped images increases performance for both races and reduces the bias, whilst adding oversampling as a bias mitigation technique with cropped images reduces the bias further.
Keywords:
Bias mitigation CMR imaging Segmentation.
1 Introduction
Artificial intelligence (AI) is increasingly being used to aid medical diagnosis, prognosis and treatment planning. However, AI models have been shown to exhibit bias by protected attributes in many different applications. For example, in cardiac magnetic resonance (CMR) image segmentation, both sex [1] and race [2, 3, 4] bias have been reported. Sex bias has also been found in chest X-ray image classification [5, 6]. AI bias can have detrimental downstream impacts in medical imaging applications. For example, CMR segmentations are used to derive biomarkers whose values impact patient management, so greater errors in these biomarkers for certain protected groups can lead to inappropriate treatment choices and worse outcomes [4].
Previous work has aimed to address biases in AI models for medical imaging tasks by using bias mitigation methods. [7] evaluated and mitigated bias in chest X-ray classification models using methods based on oversampling minority protected groups and Pareto minimax optimisation. [8] proposed a framework with eleven algorithms which aimed to measure and mitigate biases in medical imaging classification datasets. [9] evaluated stratified batch sampling, a balanced dataset model and a protected group-specific model for orthopaedic image segmentation. In CMR segmentation, [3] proposed oversampling of minority protected groups and a multi-task learning approach which learnt both a segmentation model and a protected attribute classifier.
Recently, an investigation was performed to discover the root cause(s) of AI-based CMR segmentation bias [10]. One key finding of this work was that distributional differences in areas outside the heart were significant in the learning of biased representations in AI segmentation models. This finding opens up new avenues of enquiry in bias mitigation, which we explore in this paper.
The contribution of this paper is to perform the most extensive and comprehensive investigation to date of bias mitigation methods in AI-based CMR segmentation. We investigate multiple mitigation methods as well as combinations of these methods. We also investigate AI models trained using CMR images that are cropped to remove features outside of the heart that may be leading to bias, and evaluate the effectiveness of bias mitigation techniques in this setting. Our results show that CMR segmentation bias can be mitigated to produce results that have high accuracy but are not significantly different between races.
2 Methods
We use CMR images from the UK Biobank [11]. The dataset consists of end diastolic (ED) and end systolic (ES) cine short-axis images from 5,778 subjects. Manual segmentation of the left ventricular blood pool (LVBP), left ventricular myocardium (LVM), and right ventricular blood pool (RVBP) was performed for the ED and ES images of each subject. The LV endocardial and epicardial borders and the RV endocardial border were outlined using cvi42 (version 5.1.1, Circle Cardiovascular Imaging Inc., Calgary, Alberta, Canada). The same guidelines were provided to a panel of ten experts with one expert annotating each image. Each expert was provided with a random selection of images for annotation which included subjects of different sexes and races. They were not provided with demographic information about the subjects.
Previous work [2, 1] has shown that bias is greater when the imbalance between races in the training set is greater. Therefore, we curate a dataset where biases would be significant to allow for better evaluation of mitigation methods. Our training set comprises 15 Black subjects and 4,221 White subjects.
2.1 Baseline Model
We trained a baseline 2D nnU-Net v1 model [12] to segment the LVM, LVBP and RVBP using the training parameters from [1]. The demographic information of the subjects used in the training and testing datasets can be seen in Table 1 and Table 2. Training was performed using only Black and White subjects but testing was performed using all other available subjects including Mixed, Asian, Chinese and Other.
Table 1: Characteristics of subjects used in the training dataset. Mean (standard deviation) values are presented for each characteristic. Statistically significant differences between subject groups and the overall average are indicated with an asterisk * (p $<$ 0.05) and were determined using a two-tailed Student’s t-test.
Health measure Overall White Black # subjects 4236 4221 15 Age (years 64.6 (7.7) 64.6 (7.7) 57.0 (4.9)* Weight (kg) 77.0 (15.0) 77.0 (15.0) 74.8 (11.9)* Standing height (cm) 169.5 (9.2) 169.5 (9.2) 168.8 (6.9)* Body Mass Index 26.7 (4.3) 26.7 (4.3) 26.2 (3.7)*
Table 2: Characteristics of subjects used in the test dataset. Mean (standard deviation) values are presented for each characteristic. Statistically significant differences between subject groups and the overall average are indicated with an asterisk * (p $<$ 0.05) and were determined using a two-tailed Student’s t-test.
Health measure Overall White Mixed Asian Black Chinese Other # subjects 1542 469 170 387 223 111 182 Age 61.7 (7.9) 64.8 (7.7)* 59.8 (7.2)* 61.0 (8.2)* 59.1 (7.1)* 59.7 (6.4)* 62.1 (7.5) Standing height 167.4 (9.2) 169.8 (9.4)* 166.4 (8.5) 166.7 (8.6) 168.7 (9.4) 162.4 (7.1)* 165.4 (9.7)* Weight 75.1 (15.1) 77.7 (15.1)* 75.1 (15.3) 72.3 (12.8)* 82.0 (16.0)* 63.5 (9.9)* 73.3 (15.5) Body Mass Index 26.7 (4.4) 26.9 (4.2) 27.1 (5.2) 25.9 (3.7)* 28.8 (5.1)* 24.0 (3.1)* 26.6 (4.4)
2.2 Oversampling
We also trained a 2D nnU-Net model using the same data as the baseline but applied oversampling during batch selection [13]. Oversampling refers to the process of increasing the sampling of a minority group in a dataset. Here, we oversample the Black subjects in the training dataset so that they were equal to the number of White subjects in each batch used during training. This was performed using random sampling with replacement so each subject could in principle be selected more than once in a training batch. The same training parameters from [1] were used.
2.3 Reweighing
An nnU-Net model was also trained using a reweighing mitigation strategy [13]. Reweighing refers to the process of increasing the importance of under-represented groups to the model. This can be done as a pre-processing step or during training by adapting the loss function. We performed this during training by adding a weighting term to the combined Cross Entropy (CE)-Dice loss function of the nnU-Net. Each group was weighted inversely proportionally to the group size, as shown in Eq. 2. The weights were then normalised so that they summed to 1. The training parameters from [1] were again used.
Weights per group: $\quad w_{g}=\frac{n_{G}}{n_{g}+\epsilon},\quad\text{where }\epsilon=10^{-6}.$ (1)
Normalized weights: $\quad\hat{w}_{g}=\frac{w_{g}}{\sum_{j=1}^{N_{G}}w_{j}},\quad g=1,2,\dots,N_{G}.$ (2)
where $n_{g}$ is the number of samples in protected group $g$, $n_{G}$ is the number of samples in all groups and $N_{G}$ is the number of groups.
2.4 Group Distributionally Robust Optimisation
The final mitigation approach was Group Distributionally Robust Optimisation, or Group DRO, which was first proposed in [14]. The method aims to optimise the performance of the worst-performing group in a dataset. The Group DRO loss function can be formalised as:
${L_{DRO}}=max_{g\in G}\frac{1}{{{n_{g}}}}\sum\nolimits_{i\in g}{L\left(y_{i},\hat{y_{i}}\right)}$ (2)
where $L$ is the loss function computed between predicted labels $\hat{y_{i}}$ and ground truth labels $y_{i}$.
In this method, CE loss was used instead of CE-Dice loss which was used for the oversampling and reweighing experiments. Group DRO uses losses from individual samples to calculate the average loss for the groups. However, Dice loss is calculated using global statistics of the true positives, false positives and false negatives for a group or batch. It is non-additive as the numerator and denominator will change if the calculation is performed on a per-sample basis rather than for a group or batch. For Dice loss, the average group loss and global loss are different, which causes instability in training.
The code for the bias mitigation models is available at https://github.com/tiarnaleeKCL/nnUNet-bias-mitigation
2.5 Training Using Cropped Images
Following the findings of [10], which found that areas outside the heart were a contributing factor to CMR segmentation performance bias, we also performed experiments using all of the above techniques for a nnU-Net model trained using cropped CMR images. The images were cropped around the heart using a bounding box defined based on the ground truth segmentation. All images were cropped to the same size, i.e. the size of the largest heart in the dataset.
2.6 Training Using Combinations of Mitigation Methods
The final experiment combined the mitigation methods into pairs to test whether combinations of methods would improve performance. This results in three additional methods: oversampling + Group DRO, reweighing + Group DRO, and oversampling + reweighing. These methods were applied in the same way as above. CE loss was used for experiments with Group DRO and CE-Dice loss was used for oversampling + reweighing.
2.7 Metrics
For each of the methods described above, performance was measured by finding the overall Dice similarity coefficient (DSC) for subjects in the test set. Performance was measured using the median and inter-quartile range (IQR) of the DSC. Fairness metrics are also reported: the fairness gap (FG) as defined in [1] and skewed error ratio (SER) as defined in [3]. The SER is defined as $SER=\frac{max_{g}(1-D_{g})}{min_{g}(1-D_{g})}$ where $D_{g}$ is the median DSC for protected group $g$. This measures the ratio of the errors between the median DSCs for the worst-performaing and best-performing protected groups. A more fair model will have a FG closer to 0 and a SER closer to 1.
3 Results
The results of the baseline and the three bias mitigation methods can be seen in Table 3 and in Fig. 1. Oversampling was the only method to increase performance for the Black subjects such that there was no significant difference between the median DSC scores of the Black and White subjects. The method increased the median performance for Black subjects by 0.045. Fairness performance metrics (SER and FG) also decreased, showing more equitable performance. Using oversampling also caused performance to increase for the other races compared to the baseline, as can be seen in Fig. 1(a) and Fig. 1(b). This is perhaps surprising as performance on these other races was not optimised during training.
The other mitigation methods did not significantly improve performance for the Black subjects. Reweighing resulted in worse performance for the Black subjects, with the median DSC decreasing. Group DRO resulted in increased performance for the Black subjects but performance remained significantly lower than for the White subjects. Both oversampling and Group DRO slightly decreased median DSC for White subjects (although not statistically significantly), but reweighing significantly decreased median DSC for White subjects.
Table 3: DSC values for each of the bias mitigation methods. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test. * p$<$ 0.0001. The best median DSC score for Black and White subjects is shown in bold.
Baseline * Oversampling p = 0.22 Reweighing * Group DRO * White Black White Black White Black White Black Median 0.896 0.846 0.894 0.891 0.891 0.831 0.893 0.865 IQR 0.046 0.064 0.045 0.049 0.049 0.069 0.047 0.051 SER 1.486 1.032 1.544 1.260 Fairness gap 0.050 0.003 0.059 0.028
Baseline
Oversampling
Reweighing
Group DRO
Figure 1: Overall DSC for bias mitigation methods on uncropped images. The dashed line indicates median DSC for White test subjects.
As shown in Table 4, the baseline model trained using cropped images improved performance for both White and Black subjects and reduced bias. Oversampling significantly improved performance for the Black subjects compared to the baseline, resulting in performance that was higher than for White subjects. Group DRO improved the DSC for Black subjects but this increase was not significant. Reweighing made performance slightly worse for both races, with median DSC decreasing and SER and FG measures increasing.
Table 4: DSC values for each of the bias mitigation methods using cropped images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test. * p$<$ 0.0001. The best median DSC score for Black and White subjects is shown in bold.
Baseline * Oversampling * Reweighing * Group DRO * White Black White Black White Black White Black Median 0.919 0.893 0.913 0.922 0.916 0.887 0.918 0.908 IQR 0.032 0.041 0.030 0.039 0.033 0.043 0.033 0.038 SER 1.326 1.115 1.352 1.130 Fairness gap 0.026 --0.009 0.029 0.011
Finally, combining the mitigation methods did not produce significantly less biased results, as shown in Table 5. All three combinations decreased performance for White subjects but increased performance for Black subjects. The best combination was oversampling and Group DRO which reduced the performance for the White subjects the least and improved performance for the Black subjects the most, leading to the lowest FG.
Table 5: DSC values for the combined bias mitigation methods using original sized images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test. * p$<$ 0.0001. The best median DSC score for Black and White subjects is shown in bold.
Baseline Oversampling + Group DRO Group DRO + Reweighing Oversampling + Reweighing White Black White Black White Black White Black Median 0.896 0.846 0.889 0.881 0.893 0.863 0.889 0.871 IQR 0.046 0.064 0.048 0.054 0.047 0.055 0.051 0.057 SER 1.486 1.08 1.285 1.161 Fairness gap 0.050 0.009 0.031 0.017
4 Discussion
This work has performed a comprehensive examination of bias mitigation methods for AI segmentation models used for cine CMR images. We have shown that bias in CMR segmentation models can be mitigated by using such methods. In particular, oversampling minority subjects reduces bias so that there is no significant difference between the performance of the Black and White subjects. Although oversampling did not add any extra information to the dataset, the method allowed the network to train on Black subjects more frequently than if oversampling was not used, allowing for better balance between protected groups. It could be anticipated that training using a small number of (oversampled) Black subjects would increase the risk of overfitting to those subjects, having a detrimental effect on generalisation, but this effect was not seen in our experiments as test performance remained high. Previous work in [10] showed that race could be classified from cine CMR images, indicating that there are distinct features in the images of different races that are recognisable to AI models. Using oversampling will allow the network to see more of these distinct features to learn better representations of the under-represented group.
Reweighing did not improve segmentation performance, instead decreasing performance for both protected groups. This may be due to increased importance being given to a small group of subjects, decreasing focus on the larger group. As described in [2] and [1], when White subjects comprised 75% of the training set, their segmentation performance was still lower than for the Black subjects who comprised 25% of the training set. This suggests that segmentation of the White subjects’ images may be a more difficult task as there may be more outliers and variation in the hearts than in the Black subjects. Reweighing these White subjects so that their importance is lower in the loss function could decrease accuracy.
Using cropped images increased performance and reduced bias compared to using uncropped images. Using oversampling with cropped images reduced bias further, but interestingly not to the same extent as using uncropped images. Also, note that to use a model trained using cropped images in clinical practice, a region-of-interest proposal network would be required to identify the cropping region at test time [15].
The combination of Group DRO and oversampling produced a model which improved performance for Black subjects but was still significantly lower than for White subjects. Therefore, using a combined method was better than using Group DRO alone but worse than oversampling alone.
This work has some limitations. For example, only Black and White subjects were used during training to control for other factors which may affect mitigation methods. Future work could investigate the effect of mitigating biases using multiple races or in other protected attributes such as age and socioeconomic status. It would also be useful to investigate the performance of bias mitigation strategies under external validation.
5 Acknowledgements
This work was supported by the Engineering & Physical Sciences Research Council Doctoral Training Partnership (EPSRC DTP) grant EP/T517963/1. This research has been conducted using the UK Biobank Resource under Application Number 17806

### v2

@@
-Does a Rising Tide Lift All Boats? Bias Mitigation for AI-based CMR Segmentation
+
 Abstract
 Artificial intelligence (AI) is increasingly being used for medical imaging tasks.
-However, there can be biases in the resulting models, particularly when they were trained using imbalanced training datasets.
-One such example has been the strong race bias effect in cardiac magnetic resonance (CMR) image segmentation models.
+However, there can be biases in AI models, particularly when they are trained using imbalanced training datasets.
+One such example has been the strong ethnicity bias effect in cardiac magnetic resonance (CMR) image segmentation models.
 Although this phenomenon has been reported in a number of publications, little is known about the effectiveness of bias mitigation algorithms in this domain.
 We aim to investigate the impact of common bias mitigation methods to address bias between Black and White subjects in AI-based CMR segmentation models.
-Specifically, we use oversampling, importance reweighing and Group DRO as well as combinations of these techniques to mitigate the race bias.
-Furthermore, motivated by recent findings on the root causes of AI-based CMR segmentation bias, we evaluate the same methods using models trained and evaluated on cropped CMR images.
+Specifically, we use oversampling, importance reweighing and Group DRO as well as combinations of these techniques to mitigate the ethnicity bias.
+Second, motivated by recent findings on the root causes of AI-based CMR segmentation bias, we evaluate the same methods using models trained and evaluated on cropped CMR images.
 We find that bias can be mitigated using oversampling, significantly improving performance for the underrepresented Black subjects whilst not significantly reducing the majority White subjects’ performance.
-Group DRO also improves performance for Black subjects but not significantly, while reweighing decreases performance for Black subjects.
-Using a combination of oversampling and Group DRO also improves performance for Black subjects but not significantly.
-Using cropped images increases performance for both races and reduces the bias, whilst adding oversampling as a bias mitigation technique with cropped images reduces the bias further.
-Keywords:
-Bias mitigation CMR imaging Segmentation.
+Using cropped images increases performance for both ethnicities and reduces the bias, whilst adding oversampling as a bias mitigation technique with cropped images reduces the bias further.
+When testing the models on an external clinical validation set, we find high segmentation performance and no statistically significant bias.
 1 Introduction
 Artificial intelligence (AI) is increasingly being used to aid medical diagnosis, prognosis and treatment planning.
-However, AI models have been shown to exhibit bias by protected attributes in many different applications.
-For example, in cardiac magnetic resonance (CMR) image segmentation, both sex [1] and race [2, 3, 4] bias have been reported.
-Sex bias has also been found in chest X-ray image classification [5, 6].
+However, AI models have been shown to exhibit bias by protected attributes in many different applications (Larrazabal et al., 2020; Seyyed-Kalantari et al., 2021a; Klingenberg et al., 2023), including AI-based segmentation of cardiac magnetic resonance (CMR) images (Puyol-Antón et al., 2021, 2022; Lee et al., 2022, 2023).
 AI bias can have detrimental downstream impacts in medical imaging applications.
-For example, CMR segmentations are used to derive biomarkers whose values impact patient management, so greater errors in these biomarkers for certain protected groups can lead to inappropriate treatment choices and worse outcomes [4].
-Previous work has aimed to address biases in AI models for medical imaging tasks by using bias mitigation methods.
-[7] evaluated and mitigated bias in chest X-ray classification models using methods based on oversampling minority protected groups and Pareto minimax optimisation.
-[8] proposed a framework with eleven algorithms which aimed to measure and mitigate biases in medical imaging classification datasets.
-[9] evaluated stratified batch sampling, a balanced dataset model and a protected group-specific model for orthopaedic image segmentation.
-In CMR segmentation, [3] proposed oversampling of minority protected groups and a multi-task learning approach which learnt both a segmentation model and a protected attribute classifier.
-Recently, an investigation was performed to discover the root cause(s) of AI-based CMR segmentation bias [10].
+For example, CMR segmentations are used to derive biomarkers whose values impact patient management, so greater errors in these biomarkers for certain protected groups can lead to inappropriate treatment choices and worse outcomes (Puyol-Antón et al., 2022).
+Previous work has aimed to address biases in AI models for medical imaging tasks by using generic bias mitigation methods.
+For example, Zong et al.
+(2022) proposed a framework with eleven algorithms which aimed to measure and mitigate biases in medical imaging classification datasets.
+However, the fairness gains achieved by such methods often cause reduced performance for some protected groups, a phenomenon known as the ‘fairness-accuracy trade-off’ Li and Li (2024).
+Furthermore, research has suggested that, whilst such generic bias mitigation approaches may appear to reduce bias when evaluated on internal validation sets, the fairness gains often do not hold when evaluated externally (Schrouff et al., 2022; Yang et al., 2024).
+We hypothesise that a possible reason for this lack of effectiveness is that these methods take a “blind” approach to addressing bias in AI models, i.e.
+they apply a generic bias mitigation algorithm that does not take account of the underlying cause of the bias.
+More recently, a new strand of research has emerged which attempts to understand the bias, with a view to using this knowledge to develop more informed mitigation approaches.
+For example, Glocker et al.
+(2023) analysed bias in chest X-ray classifiers, applying test-set resampling, multitask learning, and model inspection to provide insights into the way protected attributes were encoded in the AI model.
+Similarly, Olesen et al.
+(2025) used ‘slice discovery methods’ to reveal the root causes of sex bias, also in chest X-ray classifiers.
+Such methods potentially enable the development of bias mitigation techniques that are targeted at the underlying cause(s) of the bias, and could lead to improved and more robust mitigation, which will hold under the effects of domain shift to external validation sets.
+Recently, an investigation was performed to discover the root cause(s) of AI-based CMR segmentation bias (Lee et al., 2025a).
 One key finding of this work was that distributional differences in areas outside the heart were significant in the learning of biased representations in AI segmentation models.
 This finding opens up new avenues of enquiry in bias mitigation, which we explore in this paper.
-The contribution of this paper is to perform the most extensive and comprehensive investigation to date of bias mitigation methods in AI-based CMR segmentation.
-We investigate multiple mitigation methods as well as combinations of these methods.
-We also investigate AI models trained using CMR images that are cropped to remove features outside of the heart that may be leading to bias, and evaluate the effectiveness of bias mitigation techniques in this setting.
-Our results show that CMR segmentation bias can be mitigated to produce results that have high accuracy but are not significantly different between races.
-2 Methods
-We use CMR images from the UK Biobank [11].
+2 Related Works
+2.1 Bias in medical imaging
+AI bias has been observed in a wide range of medical imaging modalities and tasks.
+For example, Seyyed-Kalantari et al.
+(2021a) found bias in chest X-ray classification models in terms of sex and ethnicity.
+A CNN-based model showed a favourable bias towards males and older people, with these groups having higher true positive rates.
+Similarly, Seyyed-Kalantari et al.
+(2021b) found that younger people, females, patients under 20 years old, Black patients and Hispanic patients had higher rates of under-diagnosis.
+Larrazabal et al.
+(2020) also used chest X-ray data, and investigated the effect of training a thoracic disease classification model on datasets that were imbalanced by sex.
+They found a relationship between training set representation and performance for both sexes, and that training a model with more balanced data did not significantly decrease accuracy for the majority group but significantly improved performance for the under-represented group.
+Bias in brain magnetic resonance imaging (MRI) classification has also been reported.
+A key early work here was Petersen et al.
+(2022), who reported bias in a CNN-based model for Alzheimer’s disease (AD) detection.
+Similarly, Klingenberg et al.
+(2023) found higher accuracy in MRI-based AD detection in females than males, despite balancing the training dataset by age and sex.
+Wang et al.
+(2023) also assessed bias in brain MRI disease classification, investigating the impact of training choices on bias.
+In dermatology imaging , Abbasi-Sureshjani et al.
+(2020) found bias by age and sex when training lesion classification models.
+Daneshjou et al.
+(2022) found disparities in accuracy between lighter and darker skin tones, with the model achieving higher accuracy on the lighter skin tones.
+Fine-tuning the model on a diverse dataset closed this gap in performance and improved overall model performance so that it was equal to, or exceeded, clinician accuracy.
+Puyol-Antón et al.
+(2021) was the first paper to report bias in AI-based segmentation models.
+This work found that an nnU-Net model trained on ethnicity-imbalanced CMR data produced biased results by ethnicity.
+The segmentation performance favoured White subjects, who were in the majority in the training set.
+Subsequent work has investigated the downstream clinical impact of this bias (Puyol-Antón et al., 2022) and analysed it in more controlled experiments by both ethnicity and sex (Lee et al., 2022, 2023).
+Segmentation bias has since also been reported in brain MRI (Ioannou et al., 2022), dermatology images (Benčević et al., 2024) and orthopaedic radiographs (Siddiqui et al., 2024).
+2.2 Bias mitigation methods
+Bias mitigation methods aim to reduce the bias observed between protected groups in an AI model.
+They can be applied as pre-, in- or post-processing methods (Mehrabi et al., 2019).
+Pre-processing methods aim to transform the data before training to remove bias.
+These methods include targeted data collection and importance reweighing (which can also be performed as an in-processing method).
+Efforts to collect data from more diverse populations include a dataset of dermatology data from four African countries (Gottfrois et al., 2024), a leprosy skin imaging dataset from Brazil (Barbieri et al., 2022), brain MRI images from women with fibromyalgia in Mexico (Balducci et al., 2022) and brain tumour segmentation data from Nigeria (Adewole et al., 2024).
+In-processing methods can be used to change the objective function or apply constraints to the model during training to reduce bias.
+An example of such an approach is Group Distributionally Robust Optimisation, or Group DRO (Sagawa et al., 2020), which alters the loss function to optimise performance for the worst performing group in a dataset.
+Group DRO was one of the approaches evaluated for mitigating bias in chest X-ray classifiers in the comparative analysis by Zhang et al.
+(2022).
+Another example of a modified loss function is Pareto minimax optimisation (Martinez et al., 2020), which aims to minimise an importance-weighted maximum ‘risk’ across protected groups.
+Adversarial learning has also been suggested as an in-processing method for bias mitigation.
+In Madras et al.
+(2018), a model was adversarially trained to learn fair representations using an encoder-decoder network.
+Similarly, Zhang et al.
+(2018) used adversarial debiasing with the added constraint of satisfying fairness definitions such as demographic parity, equalised odds or equalised opportunity.
+Post-processing methods aim to modify the predictions made by the model based on a subject’s protected attributes.
+Lohia et al.
+(2019) proposed an algorithm which assesses an individual sample’s prediction, establishes whether it is biased and changes the prediction to the privileged group’s label if the sample experiences bias.
+Reject option classification, proposed in Kamiran et al.
+(2012), considers the confidence of predictions.
+For a binary classifier, prediction probabilities close to 0 or 1 represent confident predictions, whereas probabilities close to 0.5 represent more uncertain predictions.
+Samples are not assigned a label if their probabilities lie within a certain uncertainty range as they are considered more prone to bias and are relabelled depending on the group they belong to.
+Bias mitigation in AI-based segmentation has received relatively little attention compared to classification tasks.
+Relevant works here include FairSeg (Tian et al., 2023), which published a fairness dataset for medical image segmentation and proposed a fair error-bound scaling approach to reweight the loss function.
+Siddiqui et al.
+(2024) evaluated stratified batch sampling, a balanced dataset model and a protected group-specific model for orthopaedic image segmentation.
+Finally, in CMR segmentation, Puyol-Antón et al.
+(2021) internally evaluated protected group-specific models, oversampling of minority protected groups and a multi-task learning approach which learnt both a segmentation model and a protected attribute classifier.
+3 Contributions
+In the context of AI-based segmentation of cine CMR, the main contributions of this novel work are:
+  1. 1.
+    We perform the most extensive and comprehensive investigation to date of multiple bias mitigation methods in CMR segmentation, as well as combinations of these methods.
+  2. 2.
+    We perform one of the first investigations into using knowledge of the root cause of bias for mitigation.
+Specifically, we train AI segmentation models using CMR images that are cropped to remove features outside of the heart.
+We evaluate a baseline model and state-of-the-art bias mitigation techniques in this setting.
+  3. 3.
+    We evaluate the efficacy of all approaches under both internal and external validation settings.
+A preliminary version of this work has been published in Lee et al.
+(2025b).
+This paper extends that work by (i) incorporating a wider range of metrics and more in-depth analysis and discussion, (ii) inclusion of a new clinically-applicable ‘cascaded’ cropping based mitigation approach and (iii) including external validation of all approaches on a clinical dataset.
+4 Materials and Methods
+4.1 Data
+To train and internally validate all models we used CMR images from the UK Biobank (Petersen et al., 2016).
 The dataset consists of end diastolic (ED) and end systolic (ES) cine short-axis images from 5,778 subjects.
@@
 The same guidelines were provided to a panel of ten experts with one expert annotating each image.
-Each expert was provided with a random selection of images for annotation which included subjects of different sexes and races.
+Each expert was provided with a random selection of images for annotation which included subjects of different sexes and ethnicities.
 They were not provided with demographic information about the subjects.
-Previous work [2, 1] has shown that bias is greater when the imbalance between races in the training set is greater.
-Therefore, we curate a dataset where biases would be significant to allow for better evaluation of mitigation methods.
-Our training set comprises 15 Black subjects and 4,221 White subjects.
-2.1 Baseline Model
-We trained a baseline 2D nnU-Net v1 model [12] to segment the LVM, LVBP and RVBP using the training parameters from [1].
-The demographic information of the subjects used in the training and testing datasets can be seen in Table 1 and Table 2.
-Training was performed using only Black and White subjects but testing was performed using all other available subjects including Mixed, Asian, Chinese and Other.
+Previous work (Lee et al., 2022, 2023) has shown that bias is greater when the imbalance between ethnicities in the training set is greater.
+Therefore, we curated a dataset where biases would be significant to allow for better evaluation of mitigation methods.
+Our training set comprises 15 Black subjects and 4,221 White subjects, all randomly sampled from the full dataset.
+The remaining subjects were used as the internal validation test set.
+The demographic information of the subjects in the training and test sets can be seen in Table 1 and Table 2.
+Note that training was performed using only Black and White subjects but testing was performed using all other available subjects including Mixed, Asian, Chinese and Other.
+All ethnicity information refers to self-reported ethnicity based on the categories provided in the UK Biobank dataset.
+In addition, for external validation a dataset of cine short-axis CMR images from St.
+Thomas’ Hospital, London was used.
+All subjects were scanned using a 1.5T MRI scanner (Aera-Magneton, Siemens Healthcare, Erlangen, Germany) between November 2019 and November 2021.
+The patients scanned had suspected hypertension.
+Ethnicity was recorded as Black if both parents self-identified as African descendants or White if both parents self-identified as European descendants.
+The demographic information for the subjects used can be seen in Table 3.
+Ground truth segmentations were performed manually by clinical experts using cvi42 (Circle Cardiovascular Imaging Inc., Calgary, Alberta, Canada).
+Further details of the imaging protocol can be found in Georgiopoulos et al.
+(2024).
 Table 1: Characteristics of subjects used in the training dataset.
@@
 Statistically significant differences between subject groups and the overall average are indicated with an asterisk * (p $<$ 0.05) and were determined using a two-tailed Student’s t-test.
-Health measure Overall White Black # subjects 4236 4221 15 Age (years 64.6 (7.7) 64.6 (7.7) 57.0 (4.9)* Weight (kg) 77.0 (15.0) 77.0 (15.0) 74.8 (11.9)* Standing height (cm) 169.5 (9.2) 169.5 (9.2) 168.8 (6.9)* Body Mass Index 26.7 (4.3) 26.7 (4.3) 26.2 (3.7)*
-Table 2: Characteristics of subjects used in the test dataset.
+Health measure Overall White Black n subjects 4236 4221 15 Age (years) 64.6 (7.7) 64.6 (7.7) 57.0 (4.9)* Weight (kg) 77.0 (15.0) 77.0 (15.0) 74.8 (11.9)* Height (cm) 169.5 (9.2) 169.5 (9.2) 168.8 (6.9)* Body Mass Index 26.7 (4.3) 26.7 (4.3) 26.2 (3.7)
+Table 2: Characteristics of subjects used in the internal validation test dataset.
+Mean (standard deviation) values are presented for each characteristic.
+One-way ANOVA was used to test for differences across ethnic groups.
+Statistically significant differences between groups are indicated with an asterisk.
+* (p $<$ 0.01).
+Health measure Overall White Mixed Asian Black Chinese Other n subjects 1542 469 170 387 223 111 182 Age (years) * 61.7 (7.9) 64.8 (7.7) 59.8 (7.2) 61.0 (8.2) 59.1 (7.1) 59.7 (6.4) 62.1 (7.5) Height (cm) * 167.4 (9.2) 169.8 (9.4) 166.4 (8.5) 166.7 (8.6) 168.7 (9.4) 162.4 (7.1) 165.4 (9.7) Weight (kg) * 75.1 (15.1) 77.7 (15.1) 75.1 (15.3) 72.3 (12.8) 82.0 (16.0) 63.5 (9.9) 73.3 (15.5) Body Mass Index * 26.7 (4.4) 26.9 (4.2) 27.1 (5.2) 25.9 (3.7) 28.8 (5.1) 24.0 (3.1) 26.6 (4.4)
+Table 3: Characteristics of subjects used in the external validation test set.
 Mean (standard deviation) values are presented for each characteristic.
 Statistically significant differences between subject groups and the overall average are indicated with an asterisk * (p $<$ 0.05) and were determined using a two-tailed Student’s t-test.
-Health measure Overall White Mixed Asian Black Chinese Other # subjects 1542 469 170 387 223 111 182 Age 61.7 (7.9) 64.8 (7.7)* 59.8 (7.2)* 61.0 (8.2)* 59.1 (7.1)* 59.7 (6.4)* 62.1 (7.5) Standing height 167.4 (9.2) 169.8 (9.4)* 166.4 (8.5) 166.7 (8.6) 168.7 (9.4) 162.4 (7.1)* 165.4 (9.7)* Weight 75.1 (15.1) 77.7 (15.1)* 75.1 (15.3) 72.3 (12.8)* 82.0 (16.0)* 63.5 (9.9)* 73.3 (15.5) Body Mass Index 26.7 (4.4) 26.9 (4.2) 27.1 (5.2) 25.9 (3.7)* 28.8 (5.1)* 24.0 (3.1)* 26.6 (4.4)
+Health measure Overall White Black n subjects 84 30 54 Age (years) 46.0 (12.8) 40.5 (14.1) 49.1 (11.0) Weight (kg) 94.6 (19.7) 92.6 (19.3) 95.1 (20.2) Height (cm) 174.3 (8.96) 176.9 (7.33) 172.9 (9.48) Body Mass Index 31.2 (6.23) 29.6 (6.13) 31.8 (6.23)
+4.2 Baseline Model
+As a baseline model, we trained a 2D nnU-Net v1 model (Isensee et al., 2020) using the UK Biobank training set to segment the LVM, LVBP and RVBP.
 4.3 Oversampling
-We also trained a 2D nnU-Net model using the same data as the baseline but applied oversampling during batch selection [13].
+We also trained a 2D nnU-Net model using the same data as the baseline but applied oversampling during batch selection (Kamiran and Calders, 2012).
 Oversampling refers to the process of increasing the sampling of a minority group in a dataset.
-Here, we oversample the Black subjects in the training dataset so that they were equal to the number of White subjects in each batch used during training.
+Here, we oversample the Black subjects in the training set so that they were equal to the number of White subjects in each batch used during training.
 This was performed using random sampling with replacement so each subject could in principle be selected more than once in a training batch.
-The same training parameters from [1] were used.
 4.4 Reweighing
-An nnU-Net model was also trained using a reweighing mitigation strategy [13].
+An nnU-Net model was also trained using a reweighing mitigation strategy (Kamiran and Calders, 2012).
 Reweighing refers to the process of increasing the importance of under-represented groups to the model.
-This can be done as a pre-processing step or during training by adapting the loss function.
-We performed this during training by adding a weighting term to the combined Cross Entropy (CE)-Dice loss function of the nnU-Net.
+We implemented this strategy by adding a weighting term to the combined Cross Entropy (CE)-Dice loss function of the nnU-Net.
 Each group was weighted inversely proportionally to the group size, as shown in Eq.
+1.
+The weights were then normalised so that they summed to 1, as shown in Eq.
 2.
-The weights were then normalised so that they summed to 1.
-The training parameters from [1] were again used.
 Weights per group: $\quad w_{g}=\frac{n_{G}}{n_{g}+\epsilon},\quad\text{where }\epsilon=10^{-6}.$ (1)
@@
 4.5 Group Distributionally Robust Optimisation
-The final mitigation approach was Group Distributionally Robust Optimisation, or Group DRO, which was first proposed in [14].
+The final mitigation approach was Group DRO, which was first proposed in Sagawa et al.
+(2020).
 The method aims to optimise the performance of the worst-performing group in a dataset.
@@
 In this method, CE loss was used instead of CE-Dice loss which was used for the oversampling and reweighing experiments.
-Group DRO uses losses from individual samples to calculate the average loss for the groups.
+The reason for this is that Group DRO uses losses from individual samples to calculate the average loss for the groups.
 However, Dice loss is calculated using global statistics of the true positives, false positives and false negatives for a group or batch.
@@
 For Dice loss, the average group loss and global loss are different, which causes instability in training.
-The code for the bias mitigation models is available at https://github.com/tiarnaleeKCL/nnUNet-bias-mitigation
-2.5 Training Using Cropped Images
-Following the findings of [10], which found that areas outside the heart were a contributing factor to CMR segmentation performance bias, we also performed experiments using all of the above techniques for a nnU-Net model trained using cropped CMR images.
-The images were cropped around the heart using a bounding box defined based on the ground truth segmentation.
-All images were cropped to the same size, i.e.
-the size of the largest heart in the dataset.
 4.6 Training Using Combinations of Mitigation Methods
-The final experiment combined the mitigation methods into pairs to test whether combinations of methods would improve performance.
+We also combined the mitigation methods into pairs to test whether combinations of methods would improve performance.
 This results in three additional methods: oversampling + Group DRO, reweighing + Group DRO, and oversampling + reweighing.
@@
 CE loss was used for experiments with Group DRO and CE-Dice loss was used for oversampling + reweighing.
-2.7 Metrics
-For each of the methods described above, performance was measured by finding the overall Dice similarity coefficient (DSC) for subjects in the test set.
-Performance was measured using the median and inter-quartile range (IQR) of the DSC.
-Fairness metrics are also reported: the fairness gap (FG) as defined in [1] and skewed error ratio (SER) as defined in [3].
-The SER is defined as $SER=\frac{max_{g}(1-D_{g})}{min_{g}(1-D_{g})}$ where $D_{g}$ is the median DSC for protected group $g$.
-This measures the ratio of the errors between the median DSCs for the worst-performaing and best-performing protected groups.
-A more fair model will have a FG closer to 0 and a SER closer to 1.
+4.7 Training Using Cropped Images
+Following the findings of Lee et al.
+(2025a), which found that areas outside the heart were a contributing factor to CMR segmentation performance bias, we also performed experiments using all of the above techniques for a nnU-Net model trained using cropped CMR images.
+The images were cropped around the heart using a bounding box defined based on a segmentation mask.
+All images were cropped to the same size, i.e.
+the size of the largest heart in the training set plus a buffer of 5 pixels in both the $x$ and $y$ directions.
+When using cropped images, we evaluated two different approaches.
+First, we used the ground truth segmentations to calculate the cropping region at both training and inference time.
+Note that such a technique could not be used when the model is deployed since ground truth segmentations would not be available at inference time.
+Therefore, we evaluate this method to establish an upper bound on the performance of the cropping-based mitigation approach.
+Second, we developed a ‘cascaded’ cropping-based approach, in which the cropping region was estimated using an initial nnU-Net-based segmentation.
+This first nnU-Net model was trained using the full images and its output was used to estimate the cropping region.
+The resulting cropped image was then used as input to a second nnU-Net model which was trained using cropped images.
+The croppings used on the training data in this second nnU-Net were based the ground truth segmentations, which are available at training time.
+This approach is illustrated in Fig.
+1.
+Figure 1: ‘Cascaded’ approach to bias mitigation based on using cropped images.
+Images (a) are first segmented using a full-image nnU-Net (b) to produce a segmentation (c).
+This segmentation is then used to crop the images (d) and used in another nnU-Net model trained using ground truth segmentation-based cropped images (e) to produce the final segmentations (f).
+4.8 Evaluation
+For each of the methods described above, performance was measured by finding the overall Dice similarity coefficient (DSC) and Hausdorff distance (HD) for subjects in the internal and external validation test sets.
+Performance was quantified using the median and inter-quartile range (IQR) of the DSC/HD.
+Unless otherwise stated, we report the values of these metrics averaged over the LVBP, LVM and RVBP.
+Fairness metrics are also reported.
+The fairness gap (FG), represents the difference in median DSC between the protected groups (i.e.
+ethnicities) and is given by $FG=D_{White}-D_{Black}$ where $D$ is the median DSC.
+The skewed error ratio (SER), as defined in Puyol-Antón et al.
+(2021), is given by $SER=\frac{max_{g}(1-D_{g})}{min_{g}(1-D_{g})}$ where $D_{g}$ is the median DSC for protected group $g$.
+This measures the ratio of the errors between the median DSCs for the worst-performing and best-performing protected groups.
+For both fairness metrics a low value represents less bias and a more fair model will have a FG closer to 0 and a SER closer to 1.
 5 Results
-The results of the baseline and the three bias mitigation methods can be seen in Table 3 and in Fig.
-1.
+5.1 Internal validation
+The internal validation results of the baseline and bias mitigation methods can be seen in Table 4 and Fig.
+2.
 Oversampling was the only method to increase performance for the Black subjects such that there was no significant difference between the median DSC scores of the Black and White subjects.
-The method increased the median performance for Black subjects by 0.045.
+The method increased the median DSC for Black subjects by 0.045.
 Fairness performance metrics (SER and FG) also decreased, showing more equitable performance.
-Using oversampling also caused performance to increase for the other races compared to the baseline, as can be seen in Fig.
-1(a) and Fig.
-1(b).
-This is perhaps surprising as performance on these other races was not optimised during training.
+Using oversampling also caused performance to increase for the other ethnicities compared to the baseline (see Fig.
+2).
+This is perhaps surprising as performance on these other ethnicities was not optimised during training.
+Figure 2: Overall DSC for bias mitigation methods on uncropped images. The dashed line indicates median DSC for White test subjects.
 The other mitigation methods did not significantly improve performance for the Black subjects.
-Reweighing resulted in worse performance for the Black subjects, with the median DSC decreasing.
+Reweighing resulted in worse performance for the Black subjects, with the median DSC decreasing and HD increasing.
 Group DRO resulted in increased performance for the Black subjects but performance remained significantly lower than for the White subjects.
 Both oversampling and Group DRO slightly decreased median DSC for White subjects (although not statistically significantly), but reweighing significantly decreased median DSC for White subjects.
-Table 3: DSC values for each of the bias mitigation methods.
-The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test.
-* p$<$ 0.0001.
+Table 4: DSC and HD values for each of the bias mitigation methods tested on the internal validation set.
+The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
+* p$<$ 0.05.
 The best median DSC score for Black and White subjects is shown in bold.
-Baseline * Oversampling p = 0.22 Reweighing * Group DRO * White Black White Black White Black White Black Median 0.896 0.846 0.894 0.891 0.891 0.831 0.893 0.865 IQR 0.046 0.064 0.045 0.049 0.049 0.069 0.047 0.051 SER 1.486 1.032 1.544 1.260 Fairness gap 0.050 0.003 0.059 0.028
-Baseline
-Oversampling
-Reweighing
-Group DRO
-Figure 1: Overall DSC for bias mitigation methods on uncropped images. The dashed line indicates median DSC for White test subjects.
+Baseline * Oversampling p = 0.22 Reweighing * Group DRO * White Black White Black White Black White Black Median DSC 0.896 0.846 0.894 0.891 0.891 0.831 0.893 0.865 IQR DSC 0.046 0.064 0.045 0.049 0.049 0.069 0.047 0.051 Median HD (mm) 6.725 9.364 6.835 7.164 7.174 10.037 6.802 8.286 IQR HD (mm) 3.707 4.725 3.713 3.861 4.207 4.902 3.581 4.157 SER 1.486 1.032 1.544 1.260 Fairness gap 0.050 0.003 0.059 0.028
+Combining the mitigation methods did not produce significantly less biased results, as shown in Table 5.
+All three combinations decreased performance for White subjects but increased performance for Black subjects.
+The best combination was oversampling and Group DRO which reduced the performance for the White subjects the least and improved performance for the Black subjects the most, leading to the lowest FG.
+Table 5: DSC and HD values for the combined bias mitigation methods tested on the internal validation set using original sized images.
+The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
+* p$<$ 0.05.
+The best median DSC score for Black and White subjects is shown in bold.
+Baseline Oversampling + Group DRO Group DRO + Reweighing Oversampling + Reweighing White Black White Black White Black White Black Median DSC 0.896 0.846 0.889 0.881 0.893 0.863 0.889 0.871 IQR DSC 0.046 0.064 0.048 0.054 0.047 0.055 0.051 0.057 Median HD (mm) 6.725 9.364 6.985 7.740 6.831 8.417 7.364 8.654 IQR HD (mm) 3.707 4.725 3.666 4.079 3.520 4.001 4.320 4.558 SER 1.486 1.08 1.285 1.161 Fairness gap 0.050 0.009 0.031 0.017
 As shown in Table 6, the baseline model trained using cropped images improved performance for both White and Black subjects and reduced bias.
 Oversampling significantly improved performance for the Black subjects compared to the baseline, resulting in performance that was higher than for White subjects.
-Group DRO improved the DSC for Black subjects but this increase was not significant.
-Reweighing made performance slightly worse for both races, with median DSC decreasing and SER and FG measures increasing.
-Table 4: DSC values for each of the bias mitigation methods using cropped images.
-The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test.
-* p$<$ 0.0001.
+Group DRO improved the DSC for Black subjects compared to the baseline but this increase was not significant.
+However, the HD for both groups of subjects decreased.
+Reweighing made performance slightly worse for both ethnicities, with median DSC decreasing and SER and FG measures increasing.
+Table 6: DSC and HD values for each of the bias mitigation methods tested on the internal validation set using cropped images.
+The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
 The best median DSC score for Black and White subjects is shown in bold.
-Baseline * Oversampling * Reweighing * Group DRO * White Black White Black White Black White Black Median 0.919 0.893 0.913 0.922 0.916 0.887 0.918 0.908 IQR 0.032 0.041 0.030 0.039 0.033 0.043 0.033 0.038 SER 1.326 1.115 1.352 1.130 Fairness gap 0.026 --0.009 0.029 0.011
-Finally, combining the mitigation methods did not produce significantly less biased results, as shown in Table 5.
-All three combinations decreased performance for White subjects but increased performance for Black subjects.
-The best combination was oversampling and Group DRO which reduced the performance for the White subjects the least and improved performance for the Black subjects the most, leading to the lowest FG.
-Table 5: DSC values for the combined bias mitigation methods using original sized images.
-The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test.
-* p$<$ 0.0001.
+Baseline * Oversampling * Reweighing * Group DRO * White Black White Black White Black White Black Median DSC 0.917 0.888 0.906 0.914 0.915 0.877 0.917 0.903 IQR DSC 0.039 0.047 0.031 0.044 0.036 0.052 0.035 0.046 Median HD (mm) 5.713 7.895 6.394 6.479 5.987 8.945 5.700 7.170 IQR HD (mm) 2.669 3.960 3.359 3.211 3.073 4.097 2.805 3.584 SER 1.340 1.088 1.448 1.161 Fairness gap 0.028 -0.008 0.038 0.013
+Table 7 and Figs.
+2(e) to 2(f) show the results for the cascaded cropping on the internal validation dataset.
+Surprisingly, using the cascaded cropping approach significantly improved the DSC score compared to using the ground-truth cropped results shown in Table 6 for all bias mitigation methods.
+We discuss a possible reason for this in Section 6.1.
+Using oversampling on the cascaded cropping approach did not improve DSC score but decreased the HD, SER and FG compared to the baseline.
+Reweighing also improved the SER and FG.
+Table 7: DSC and HD values for each of the bias mitigation methods using the cascaded cropping approach on the internal validation data.
+The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
+* p$<$ 0.05.
 The best median DSC score for Black and White subjects is shown in bold.
-Baseline Oversampling + Group DRO Group DRO + Reweighing Oversampling + Reweighing White Black White Black White Black White Black Median 0.896 0.846 0.889 0.881 0.893 0.863 0.889 0.871 IQR 0.046 0.064 0.048 0.054 0.047 0.055 0.051 0.057 SER 1.486 1.08 1.285 1.161 Fairness gap 0.050 0.009 0.031 0.017
+Baseline * Oversampling * Reweighing * Group DRO * White Black White Black White Black White Black Median DSC 0.959 0.937 0.958 0.941 0.951 0.935 0.955 0.943 IQR DSC 0.060 0.057 0.064 0.058 0.067 0.070 0.068 0.058 Median HD (mm) 3.757 6.564 3.731 4.929 5.176 6.426 3.841 4.623 IQR HD (mm) 2.268 5.012 2.129 2.717 3.262 4.172 2.039 3.327 SER 1.517 1.380 1.332 1.276 Fairness gap 0.021 0.016 0.016 0.012
+5.2 External validation
+Table 8 shows the external validation results.
+Tables S1 to S3 show the results broken down by cardiac region.
+As oversampling was found to be the best bias mitigation on internal validation, only the results for this model and the baseline model are shown.
+Note that, as the LVM is not routinely segmented for the ES frame in clinical workflows at our institution, results for the external validation dataset are shown for the ED frames only.
+Table 8: DSC and HD values for the baseline and oversampling methods tested on the external validation set.
+The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
+* p$<$ 0.05.
+The best median DSC score for Black and White subjects is shown in bold.
+Baseline p = 0.093 Oversampling p = 0.16 White Black White Black Median DSC 0.886 0.879 0.885 0.872 IQR DSC 0.030 0.043 0.037 0.038 Median HD (mm) 5.728 5.913 6.152 5.919 IQR HD (mm) 1.522 1.966 1.238 1.840 SER 1.064 1.106 Fairness gap 0.007 0.012
+Overall, the model had high performance on the external validation set.
+The performance was comparable to that of the bias mitigation methods tested on the internal validation set, with the baseline performance for the Black subjects being higher than for the Black subjects in the internal validation set.
+The fairness gap and SER were also smaller than for the internal validation set.
+However, for the external validation set, using oversampling decreased performance for both groups, with the FG, SER and median HD increasing.
+The approach based on cropping using ground truth segmentations produced a significantly better baseline performance in the external validation set for both groups compared to not using cropping, as shown in Table 9.
+Using oversampling slightly decreased the performance for both groups and resulted in a significantly different performance for Black and White subjects.
+The results from using the cascaded cropping based approach can be seen in Table 9.
+Tables S4 to S6 show the results broken down by cardiac region for the cascaded approach.
+Using this approach, performance was lower than the upper bound of using ground truth segmentations for cropping for both groups but performance remained higher than for the baseline model using uncropped images seen in Table 8.
+Performance increased for both groups when using oversampling and the FG decreased.
+Table S7 shows the results for the oversampling cascaded model broken down by sex and age.
+Table 9: DSC and HD values for the models using cropping on the external validation dataset.
+The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
+* p$<$ 0.05.
+The best median DSC score for Black and White subjects is shown in bold.
+Cropped baseline ground-truth model p = 0.069 Cropped oversampling ground-truth model Cropped baseline cascaded model p = 0.12 Cropped oversampling cascaded model p = 0.68 White Black White Black White Black White Black Median DSC 0.931 0.916 0.918 0.906 0.917 0.903 0.928 0.920 IQR DSC 0.043 0.037 0.040 0.033 0.045 0.038 0.039 0.026 Median HD (mm) 4.087 4.090 4.793 4.594 4.772 4.954 4.658 3.875 IQR HD (mm) 1.630 1.178 1.321 1.529 3.124 1.919 2.69 1.
+432 SER 1.207 1.141 1.168 1.118 Fairness gap 0.014 0.012 0.014 0.008
 6 Discussion
+6.1 Internal validation
 This work has performed a comprehensive examination of bias mitigation methods for AI segmentation models used for cine CMR images.
@@
 It could be anticipated that training using a small number of (oversampled) Black subjects would increase the risk of overfitting to those subjects, having a detrimental effect on generalisation, but this effect was not seen in our experiments as test performance remained high.
-Previous work in [10] showed that race could be classified from cine CMR images, indicating that there are distinct features in the images of different races that are recognisable to AI models.
-Using oversampling will allow the network to see more of these distinct features to learn better representations of the under-represented group.
+Previous work in Lee et al.
+(2025a) has showed that ethnicity can be classified from cine CMR images, indicating that there are distinct features in the images of different ethnicities that are recognisable to AI models.
+Using oversampling will allow the network to see more of these distinct features to learn better representations for the under-represented group.
 Reweighing did not improve segmentation performance, instead decreasing performance for both protected groups.
 This may be due to increased importance being given to a small group of subjects, decreasing focus on the larger group.
-As described in [2] and [1], when White subjects comprised 75% of the training set, their segmentation performance was still lower than for the Black subjects who comprised 25% of the training set.
+As described in Lee et al.
+(2022) and Lee et al.
+(2023), when White subjects comprised 75% of the training set, their segmentation performance was still lower than for the Black subjects who comprised 25% of the training set.
 This suggests that segmentation of the White subjects’ images may be a more difficult task as there may be more outliers and variation in the hearts than in the Black subjects.
 Reweighing these White subjects so that their importance is lower in the loss function could decrease accuracy.
+The combination of Group DRO and oversampling produced a model which improved performance for Black subjects but was still significantly lower than for White subjects.
+Therefore, using a combined method was better than using Group DRO alone but worse than oversampling alone.
 Using cropped images increased performance and reduced bias compared to using uncropped images.
 Using oversampling with cropped images reduced bias further, but interestingly not to the same extent as using uncropped images.
-Also, note that to use a model trained using cropped images in clinical practice, a region-of-interest proposal network would be required to identify the cropping region at test time [15].
-The combination of Group DRO and oversampling produced a model which improved performance for Black subjects but was still significantly lower than for White subjects.
-Therefore, using a combined method was better than using Group DRO alone but worse than oversampling alone.
+The cascaded cropping approach resulted in a performance that was higher than using ground truth segmentations for cropping.
+This may be due to the intermediate predictions on the uncropped images (c in Fig.
+1) being less accurate and the segmented area being larger than the ground truth segmentations, resulting in a bounding box which was larger and contained more information.
+This suggests that performance could be optimised further by increasing the size of the pixel buffer used in the cropping process.
+6.2 External validation
+Using the bias mitigation models on the external validation set produced comparable DSC scores to internal validation results.
+No significant bias was observed in the baseline model for this dataset, as shown in Table 8.
+This is surprising as the dataset is out-of-distribution and some distributional shift is expected, which could lead to worse performance.
+However, unlike the models tested on the internal validation set, oversampling did not decrease the FG or improve the performance of the models under external validation.
+This is consistent with previous work that has reported limited effectiveness of generic bias mitigation algorithms under complex domain shifts (Schrouff et al., 2022; Yang et al., 2024).
+Cropping the images increased performance overall, both using the ground truth-based cropping and the cascaded approach.
+Importantly, using the cascaded approach, performance was higher than the baseline model using uncropped images for both groups under external validation.
+Interestingly, using oversampling in combination with the cascaded approach improved performance for both groups further.
+These results indicate that, not only does cropping improve the performance of models for both protected groups (i.e.
+ethnicities), but also that ground truth segmentations are not needed to crop these images.
+This finding can be vital for the clinical translation of bias mitigation algorithms in CMR analysis, as better segmentation performance will allow for better assessment of clinical biomarkers and better treatment planning, prognosis and diagnosis.
+Provided that there is both a network trained on uncropped images and a network trained on cropped images, the method will be deployable and scalable.
+In the future, the first model used to localise the cardiac region of interest (b in Fig.
+1) could be replaced with a similar network trained for bounding box detection such as that proposed in He et al.
+(2017).
+The best method overall (cascaded cropping used with oversampling) was based on an understanding of the root cause of bias in AI-based CMR analysis as reported in Lee et al.
+(2025a).
+This method resulted in not only reduced bias compared to the baseline model but also improved performance for both groups.
+This important result highlights that, when using a bias understanding-informed mitigation approach, there was no fairness-accuracy trade-off.
+The fairness-accuracy trade-off is sometimes used as a reason not to implement bias mitigation algorithms in clinical practice, due to the medical principle of non-maleficence.
+Our work suggests that this trade-off is not inevitable, but that it can be avoided by a careful analysis of bias in individual scenarios rather than application of a generic mitigation algorithm.
+6.3 Limitations and future work
 This work has some limitations.
-For example, only Black and White subjects were used during training to control for other factors which may affect mitigation methods.
-Future work could investigate the effect of mitigating biases using multiple races or in other protected attributes such as age and socioeconomic status.
-It would also be useful to investigate the performance of bias mitigation strategies under external validation.
-5 Acknowledgements
-This work was supported by the Engineering & Physical Sciences Research Council Doctoral Training Partnership (EPSRC DTP) grant EP/T517963/1.
-This research has been conducted using the UK Biobank Resource under Application Number 17806
+For example, only Black and White subjects were used during training.
+Future work could investigate the effect of mitigating biases using multiple ethnicities, in other protected attributes such as age and socioeconomic status, or in intersectional groups.
+There was also a relatively small number of subjects in the external dataset when compared to the number of subjects in the training and internal validation sets from the UK Biobank, so externally validating on a larger, more diverse dataset would be beneficial.
+7 Conclusions
+This paper has reported the most comprehensive investigation of bias mitigation in AI-based CMR segmentation to date.
+We have shown that the fairness-accuracy trade-off can be avoided by using a bias understanding-informed approach to mitigation, rather than using a generic mitigation algorithm.
+Therefore, this represents an important finding that should motivate further investigations into such bias understanding-informed approaches to mitigation in other applications.
+[…]
+Engineering & Physical Sciences Research Council Doctoral Training Partnership (EPSRC DTP) grant EP/T517963/1.
+This research has been conducted using the UK Biobank Resource under Application Number 17806.
+[…]
+The work follows appropriate ethical standards in conducting research and writing the manuscript, following all applicable laws and regulations regarding treatment of animals and human subjects
+[…]
+E.P-A.
+is an employee of Heartflow Inc.
+All work presented is independent of her role at Heartfow.
+P-G.M.
+worked as a consultant for Perspectum Diagnostics Ltd until 2023.
+The remaining authors declare no conflict of interest.
+[…]
+The UK Biobank dataset is publicly available for approved research projects.
+Requests to access the dataset should be directed to https://www.ukbiobank.ac.uk/.
+The clinical dataset cannot be publicly shared due to limitations of the ethical approval.
+Supplementary Material
+Table S1: Metrics for LVBP segmentation based on inference performed on external validation set using different models trained on UK Biobank data.
+The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
+* p$<$ 0.05.
+The best median DSC score for Black and White subjects is shown in bold.
+Baseline p = 0.11 Oversampling p = 0.46 White Black White Black Median DSC 0.940 0.931 0.936 0.937 IQR DSC 0.041 0.035 0.038 0.032 Median HD (mm) 3.929 3.416 3.705 3.518 IQR HD (mm) 1.622 1.498 1.296 2.186 SER 1.146 1.007 Fairness gap 0.009 -0.0004
+Table S2: Metrics for LVM segmentation based on inference performed on the external validation set using different models trained on UK Biobank data.
+The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
+* p$<$ 0.05.
+The best median DSC score for Black and White subjects is shown in bold.
+Baseline p = 0.84 Oversampling p = 0.57 White Black White Black Median DSC 0.817 0.803 0.808 0.802 IQR DSC 0.046 0.075 0.053 0.06 Median HD (mm) 5.026 5.141 5.096 5.318 IQR HD (mm) 1.583 1.941 1.995 2.042 SER 1.077 1.034 Fairness gap 0.014 0.006
+Table S3: Metrics for RVBP segmentation based on inference performed on the external validation set using different models trained on UK Biobank data.
+The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
+* p$<$ 0.05.
+The best median DSC score for Black and White subjects is shown in bold.
+Baseline p = 0.43 Oversampling p = 0.27 White Black White Black Median DSC 0.883 0.882 0.885 0.868 IQR DSC 0.057 0.048 0.070 0.071 Median HD (mm) 8.902 7.783 8.869 10.181 IQR HD (mm) 3.208 3.096 3.962 4.401 SER 1.004 1.148 Fairness gap 0.0004 0.017
+Table S4: Metrics for LVBP segmentation based on inference performed on the external validation set using the cascaded approach to crop the images.
+The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
+* p$<$ 0.05.
+The best median DSC score for Black and White subjects is shown in bold.
+Cropped baseline p = 0.088 Cropped oversampling p = 0.76 Cropped baseline cascaded model Cropped oversampling cascaded model p = 0.72 White Black White Black White Black White Black Median DSC 0.962 0.949 0.958 0.943 0.959 0.943 0.959 0.946 IQR DSC 0.030 0.031 0.024 0.027 0.031 0.029 0.036 0.024 Median HD (mm) 2.887 2.488 3.043 2.870 3.132 2.839 2.842 2.493 IQR HD (mm) 0.984 0.854 1.113 1.697 1.119 1.049 1.352 2.646 SER 1.309 1.364 1.398 1.316 Fairness gap 0.012 0.015 0.016 0.013
+Table S5: Metrics for LVM segmentation based on inference performed on external validation set using the cascaded approach to crop the images.
+The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
+* p$<$ 0.05.
+The best median DSC score for Black and White subjects is shown in bold.
+Cropped baseline p = 0.39 Cropped oversampling p = 0.63 Cropped baseline cascaded model p = 0.69 Cropped oversampling cascaded model p = 0.86 White Black White Black White Black White Black Median DSC 0.881 0.860 0.852 0.849 0.856 0.855 0.879 0.863 IQR DSC 7 0.103 0.093 0.096 0.083 0.093 0.106 0.071 0.066 Median HD (mm) 3.512 3.310 4.071 3.937 4.051 4.008 3.422 3.259 IQR HD (mm) 1.738 1.251 2.030 2.492 2.532 2.098 2.051 3.120 SER 1.175 1.019 1.009 1.130 Fairness gap 0.021 0.003 0.001 0.016
+Table S6: Metrics for RVBP segmentation based on inference performed on external validation set using the cascaded approach to crop the images.
+The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
+* p$<$ 0.05.
+The best median DSC score for Black and White subjects is shown in bold.
+Cropped baseline p = 0.75 Cropped oversampling p = 0.71 Cropped baseline cascaded model p =0.84 Cropped oversampling cascaded model White Black White Black White Black White Black Median DSC 0.938 0.925 0.918 0.907 0.912 0.905 0.929 0.932 IQR DSC 0.024 0.022 0.034 0.026 0.045 0.040 0.049 0.031 Median HD (mm) 5.932 5.787 6.540 6.662 7.477 7.147 4.871 6.426 IQR HD (mm) 2.580 3.557 2.991 2.753 4.459 5.708 2.018 4.328 SER 1.197 1.136 1.131 1.036 Fairness gap 0.012 0.011 0.011 -0.003
+Table S7: DSC and HD values by sex and age for the cascaded approach using oversampling on the external validation set.
+The p-values were computed based on a two-sided Mann Whitney U test on the DSC scores.
+* p$<$ 0.05.
+FG was calculated using $DSC_{male}-DSC_{female}$ and $DSC_{Age<50}-DSC_{Age>50}$
+Cropped baseline cascaded model p = 0.12 Cropped oversampling cascaded model p = 0.68 Female Male Age $<$ 50 Age $>$ 50 n 28 56 52 32 Median DSC 0.913 0.927 0.922 0.922 IQR DSC 0.032 0.036 0.032 0.031 Median HD (mm) 4.069 4.143 3.869 4.208 IQR HD (mm) 1.927 1.402 1.558 1.738 SER 1.193 1.004 Fairness gap 0.0141 -0.0003

### v3

@@
 (2022) proposed a framework with eleven algorithms which aimed to measure and mitigate biases in medical imaging classification datasets.
-However, the fairness gains achieved by such methods often cause reduced performance for some protected groups, a phenomenon known as the ‘fairness-accuracy trade-off’ Li and Li (2024).
+However, the fairness gains achieved by such methods often cause reduced performance for some protected groups, a phenomenon known as the ‘fairness-accuracy trade-off’ (Li and Li, 2024).
 Furthermore, research has suggested that, whilst such generic bias mitigation approaches may appear to reduce bias when evaluated on internal validation sets, the fairness gains often do not hold when evaluated externally (Schrouff et al., 2022; Yang et al., 2024).
@@
 Therefore, we curated a dataset where biases would be significant to allow for better evaluation of mitigation methods.
-Our training set comprises 15 Black subjects and 4,221 White subjects, all randomly sampled from the full dataset.
+Our main training set comprises 15 Black subjects and 4,221 White subjects, all randomly sampled from the full dataset.
 The remaining subjects were used as the internal validation test set.
+We also investigate the effect of using different proportions of Black and White subjects in the training set in the Supplementary Material.
 The demographic information of the subjects in the training and test sets can be seen in Table 1 and Table 2.
@@
 Mean (standard deviation) values are presented for each characteristic.
-Statistically significant differences between subject groups and the overall average are indicated with an asterisk * (p $<$ 0.05) and were determined using a two-tailed Student’s t-test.
+Statistically significant differences between subject groups and the overall average are indicated with an asterisk * (p << 0.05) and were determined using a two-tailed Student’s t-test.
 Health measure Overall White Black n subjects 4236 4221 15 Age (years) 64.6 (7.7) 64.6 (7.7) 57.0 (4.9)* Weight (kg) 77.0 (15.0) 77.0 (15.0) 74.8 (11.9)* Height (cm) 169.5 (9.2) 169.5 (9.2) 168.8 (6.9)* Body Mass Index 26.7 (4.3) 26.7 (4.3) 26.2 (3.7)
@@
 Statistically significant differences between groups are indicated with an asterisk.
-* (p $<$ 0.01).
+* (p << 0.05).
 Health measure Overall White Mixed Asian Black Chinese Other n subjects 1542 469 170 387 223 111 182 Age (years) * 61.7 (7.9) 64.8 (7.7) 59.8 (7.2) 61.0 (8.2) 59.1 (7.1) 59.7 (6.4) 62.1 (7.5) Height (cm) * 167.4 (9.2) 169.8 (9.4) 166.4 (8.5) 166.7 (8.6) 168.7 (9.4) 162.4 (7.1) 165.4 (9.7) Weight (kg) * 75.1 (15.1) 77.7 (15.1) 75.1 (15.3) 72.3 (12.8) 82.0 (16.0) 63.5 (9.9) 73.3 (15.5) Body Mass Index * 26.7 (4.4) 26.9 (4.2) 27.1 (5.2) 25.9 (3.7) 28.8 (5.1) 24.0 (3.1) 26.6 (4.4)
@@
 Mean (standard deviation) values are presented for each characteristic.
-Statistically significant differences between subject groups and the overall average are indicated with an asterisk * (p $<$ 0.05) and were determined using a two-tailed Student’s t-test.
+Statistically significant differences between subject groups and the overall average are indicated with an asterisk * (p << 0.05) and were determined using a two-tailed Student’s t-test.
 Health measure Overall White Black n subjects 84 30 54 Age (years) 46.0 (12.8) 40.5 (14.1) 49.1 (11.0) Weight (kg) 94.6 (19.7) 92.6 (19.3) 95.1 (20.2) Height (cm) 174.3 (8.96) 176.9 (7.33) 172.9 (9.48) Body Mass Index 31.2 (6.23) 29.6 (6.13) 31.8 (6.23)
@@
 For both fairness metrics a low value represents less bias and a more fair model will have a FG closer to 0 and a SER closer to 1.
+4.9 Code availability
+The code to reproduce this work is available at https://github.com/tiarnaleeKCL/nnUNet-bias-mitigation.
 5 Results
@@
 2.
+Statistical tests were performed using Mann-Whitney U tests on the overall DSC scores.
 Oversampling was the only method to increase performance for the Black subjects such that there was no significant difference between the median DSC scores of the Black and White subjects.
@@
 This is perhaps surprising as performance on these other ethnicities was not optimised during training.
+Further results on the effect of different levels of oversampling, both without and without cascaded cropping, can be seen in Figure S2.
+Plots of performance (median DSC) against FG for these two approaches can be seen in Figure S3.
+The effect of using different proportions of Black and White subjects in the training set are reported in Figure S1.
+Additional experiments using Asian and White subjects can be seen in Table S9 .
 Figure 2: Overall DSC for bias mitigation methods on uncropped images. The dashed line indicates median DSC for White test subjects.
@@
 The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
-* p$<$ 0.05.
+* p<< 0.05.
 The best median DSC score for Black and White subjects is shown in bold.
@@
 Combining the mitigation methods did not produce significantly less biased results, as shown in Table 5.
+Mann-Whitney U tests were performed to test the significance of differences between the overall DSC scores.
 All three combinations decreased performance for White subjects but increased performance for Black subjects.
@@
 The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
-* p$<$ 0.05.
+* p<< 0.05.
 The best median DSC score for Black and White subjects is shown in bold.
@@
 The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
-* p$<$ 0.05.
+* p<< 0.05.
 The best median DSC score for Black and White subjects is shown in bold.
@@
 The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
-* p$<$ 0.05.
+* p<< 0.05.
 The best median DSC score for Black and White subjects is shown in bold.
@@
 The results from using the cascaded cropping based approach can be seen in Table 9.
-Tables S4 to S6 show the results broken down by cardiac region for the cascaded approach.
+Mann-Whitney U tests were used to compare the overall DSC scores.
+Furthermore, the errors in the sizes of the bounding boxes using the cascaded approach compared to those calculated using the ground-truth segmentations can be seen in Table S4.
+Tables S5 to S7 show the results broken down by cardiac region for the cascaded approach.
 Using this approach, performance was lower than the upper bound of using ground truth segmentations for cropping for both groups but performance remained higher than for the baseline model using uncropped images seen in Table 8.
@@
 The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
-* p$<$ 0.05.
+* p<< 0.05.
 The best median DSC score for Black and White subjects is shown in bold.
@@
 (2025a).
-This method resulted in not only reduced bias compared to the baseline model but also improved performance for both groups.
-This important result highlights that, when using a bias understanding-informed mitigation approach, there was no fairness-accuracy trade-off.
+Specifically, in Lee et al.
+(2025a) it was reported that the main source of the distributional shift between ethnicities (and hence the bias in segmentation performance) was outside the heart.
+The cascaded cropping approach is a simple but elegant approach to removing the source of the bias in a robust way but maintaining (and even improving) segmentation performance.
+This important result highlights that, when using a bias understanding-informed mitigation approach in the context of CMR segmentation, there is no fairness-accuracy trade-off.
 The fairness-accuracy trade-off is sometimes used as a reason not to implement bias mitigation algorithms in clinical practice, due to the medical principle of non-maleficence.
@@
 This work has some limitations.
-For example, only Black and White subjects were used during training.
+For example, only two ethnicities (Black/White and Asian/White) were used during training in our experiments.
 Future work could investigate the effect of mitigating biases using multiple ethnicities, in other protected attributes such as age and socioeconomic status, or in intersectional groups.
 There was also a relatively small number of subjects in the external dataset when compared to the number of subjects in the training and internal validation sets from the UK Biobank, so externally validating on a larger, more diverse dataset would be beneficial.
+In future work we also plan to investigate more advanced baseline bias mitigation methods than the ones used in this work.
+However, it should be noted that the literature on bias mitigation in segmentation (e.g.
+Puyol-Antón et al.
+(2021); Benčević et al.
+(2024); Siddiqui et al.
+(2024)) is limited compared to the more extensive literature in classification problems.
+(R3.1) Finally, the different loss functions used for Group DRO compared to reweighing and oversampling may have resulted in different regularisation for the models so future work could investigate the use of a single loss function.
 7 Conclusions
@@
 Supplementary Material
+Metrics for regions of the heart using the external validation set
 Table S1: Metrics for LVBP segmentation based on inference performed on external validation set using different models trained on UK Biobank data.
 The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
-* p$<$ 0.05.
+* p<< 0.05.
 The best median DSC score for Black and White subjects is shown in bold.
@@
 The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
-* p$<$ 0.05.
+* p<< 0.05.
 The best median DSC score for Black and White subjects is shown in bold.
@@
 The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
-* p$<$ 0.05.
+* p<< 0.05.
 The best median DSC score for Black and White subjects is shown in bold.
 Baseline p = 0.43 Oversampling p = 0.27 White Black White Black Median DSC 0.883 0.882 0.885 0.868 IQR DSC 0.057 0.048 0.070 0.071 Median HD (mm) 8.902 7.783 8.869 10.181 IQR HD (mm) 3.208 3.096 3.962 4.401 SER 1.004 1.148 Fairness gap 0.0004 0.017
+Difference in the size of the bounding box between images
+Table S4: Difference in the size of the bounding box between images cropped using the ground truth and cascaded approach for the internal set.
+Experiment Error $x$ (%) Error $y$ (%)
+Baseline 3.70 6.57
+Oversampling 0.0 7.30
+Reweighing 1.48 8.76
+Group DRO 0.0 0.730
+Metrics for regions of the heart for the external validation set using the cascaded cropping method
 Table S5: Metrics for LVBP segmentation based on inference performed on the external validation set using the cascaded approach to crop the images.
 The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
-* p$<$ 0.05.
+* p<< 0.05.
 The best median DSC score for Black and White subjects is shown in bold.
@@
 The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
-* p$<$ 0.05.
+* p<< 0.05.
 The best median DSC score for Black and White subjects is shown in bold.
@@
 The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores.
-* p$<$ 0.05.
+* p<< 0.05.
 The best median DSC score for Black and White subjects is shown in bold.
 Cropped baseline p = 0.75 Cropped oversampling p = 0.71 Cropped baseline cascaded model p =0.84 Cropped oversampling cascaded model White Black White Black White Black White Black Median DSC 0.938 0.925 0.918 0.907 0.912 0.905 0.929 0.932 IQR DSC 0.024 0.022 0.034 0.026 0.045 0.040 0.049 0.031 Median HD (mm) 5.932 5.787 6.540 6.662 7.477 7.147 4.871 6.426 IQR HD (mm) 2.580 3.557 2.991 2.753 4.459 5.708 2.018 4.328 SER 1.197 1.136 1.131 1.036 Fairness gap 0.012 0.011 0.011 -0.003
+DSC and HD values by sex and age for the cascaded approach
 Table S8: DSC and HD values by sex and age for the cascaded approach using oversampling on the external validation set.
 The p-values were computed based on a two-sided Mann Whitney U test on the DSC scores.
-* p$<$ 0.05.
+* p<< 0.05.
 FG was calculated using $DSC_{male}-DSC_{female}$ and $DSC_{Age<50}-DSC_{Age>50}$
-Cropped baseline cascaded model p = 0.12 Cropped oversampling cascaded model p = 0.68 Female Male Age $<$ 50 Age $>$ 50 n 28 56 52 32 Median DSC 0.913 0.927 0.922 0.922 IQR DSC 0.032 0.036 0.032 0.031 Median HD (mm) 4.069 4.143 3.869 4.208 IQR HD (mm) 1.927 1.402 1.558 1.738 SER 1.193 1.004 Fairness gap 0.0141 -0.0003
+Cropped baseline cascaded model p = 0.12 Cropped baseline cascaded model p = 0.68 Female Male Age << 50 Age >> 50 n 28 56 52 32 Median DSC 0.913 0.927 0.922 0.922 IQR DSC 0.032 0.036 0.032 0.031 Median HD (mm) 4.069 4.143 3.869 4.208 IQR HD (mm) 1.927 1.402 1.558 1.738 SER 1.193 1.004 Fairness gap 0.0141 -0.0003
+Metrics for the bias mitigation methods using White and Asian subjects
+Table S9: DSC and HD values for each of the bias mitigation methods tested on the internal validation set.
+The p-values were computed between White and Asian subjects based on a two-sided Mann Whitney U test on the DSC scores.
+* p<< 0.05.
+The best median DSC score for Asian and White subjects is shown in bold.
+Baseline * Oversampling p = 0.17 Reweighing * Group DRO * Cropped baseline cascaded model Cropped oversampling cascaded model White Asian White Asian White Asian White Asian White Asian White Asian Median DSC 0.901 0.851 0.8990 0.9039 0.8971 0.8349 0.8971 0.8704 0.9584 0.9362 0.9576 0.8949 IQR DSC 0.0452 0.0635 0.0445 0.0678 0.0501 0.0686 0.0471 0.0584 0.0559 0.0502 0.0581 0.0583 Median HD (mm) 6.761 9.585 6.783 7.204 7.051 10.057 6.688 8.628 3.957 7.254 3.862 8.093 IQR HD (mm) 3.690 4.935 3.715 3.662 4.140 5.076 3.744 4.310 2.470 5.115 2.321 5.491 SER 1.498 1.051 1.604 1.260 1.533 2.477 Fairness gap 0.0494 -0.0049 0.0621 0.0267 0.0222 0.0627
+The effect of changing the number of Black subjects in the training dataset
+Figure S1: The effect of changing the number of Black subjects in the training dataset.
+The effect of changing the level of oversampling in the training dataset
+Figure S2: The effect of changing level of oversampling of Black subjects in the training dataset.
+Median DSC against fairness gap for different levels of oversampling
+Figure S3: Median DSC against fairness gap for different levels of oversampling in experiments using oversampling and cascaded cropping + oversampling.
+The fairness gap is calculate by subtracting the median DSC for Black subjects from the median DSC for White subjects.